OpenStudy (anonymous):

solve. 1/2x-8 - 3/x-4 = -5 x=3 x=9/2 x=6 x=-24/5

OpenStudy (anonymous):

@paki @YanaSidlinskiy

OpenStudy (anonymous):

hmm hmm okay you can write this as 1/2(x-4) -3/x-4 =-5

OpenStudy (anonymous):

so 1/(x-4) *(1/2 -3)= -5

OpenStudy (anonymous):

choices are x=3 x=9/2 x=6 x=-24/5

OpenStudy (anonymous):

>_<

OpenStudy (anonymous):

do you want it solved or direct answer

OpenStudy (anonymous):

sigh

OpenStudy (anonymous):

i just wanna correct answer im kind of in a rush rn

OpenStudy (anonymous):

pfffft 6

OpenStudy (anonymous):

are u sure

OpenStudy (anonymous):

nope you are gonna fail in ur practice test

OpenStudy (anonymous):

practice test?

OpenStudy (yanasidlinskiy):

@aajugdar , Please review th Code of Conduct. Direct Answers are not allowed. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

everyone is sending me the same link tell it to that kid x(

OpenStudy (anonymous):

i didnt ask for direct answer u suggested it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

great

OpenStudy (anonymous):

so where was i

OpenStudy (anonymous):

1/(x-4) *(1/2 -3)= -5

OpenStudy (anonymous):

this can be written as 1/(x-4) *(-5)/2 =-5 so 1/x-4 = 2 so 2x-8= 1 so 2x = 9 so x = 9/2

OpenStudy (anonymous):

oh dont read the last part *code of conduct*

OpenStudy (anonymous):

thanks